OpenStudy (anonymous):

Solutions to (x+3)^2=10

OpenStudy (anonymous):

-3+-sqrt(10)

OpenStudy (anonymous):

x^2 + 9 + 6x = 10 x^2 + 6x - 1 = 0

OpenStudy (anonymous):

x^2+6x-1=0 den solve by quadratic eq